# Lavida-O: Elastic Masked Diffusion Models for Unified Multimodal Understanding and Generation

### v1

@@ Abstract
We proposed Lavida-O, a unified multi-modal Masked Diffusion Model (MDM) capable of image understanding and generation tasks. Unlike existing multimodal diffsion language models such as MMaDa and Muddit which only support simple image-level understanding tasks and low-resolution image generation, Lavida-O exhibits many new capabilities such as object grounding, image-editing, and high-resolution (1024px) image synthesis. It is also the first unified MDM that uses its understanding capabilities to improve image generation and editing results through planning and iterative self-reflection. To allow effective and efficient training and sampling, Lavida-O introduces many novel techniques such as Elastic Mixture-of-Transformer architecture, universal text conditioning, and stratified sampling. Lavida-O achieves state-of-the-art performance on a wide range of benchmarks such as RefCOCO object grounding, GenEval text-to-image generation, and ImgEdit image editing, outperforming existing autoregressive and continuous diffusion models such as Qwen2.5-VL and FluxKontext-dev, while offering considerable speedup at inference.
@@ 1 Introduction
Recently, Masked Diffusion Models (MDMs) (Lou et al., 2023; Sahoo et al., 2024) have emerged as a competitive alternative to AR models. Unlike AR models, MDMs treat token generation as a diffusion process over discrete tokens. In the forward process, the tokens of a sequence are gradually masked. At inference, we start with a sequence of mask tokens and gradually unmask them to obtain a sequence of meaningful tokens. Large-scale experiments in language modeling (Nie et al., 2025; Ye et al., 2025a) show that MDMs can achieve comparable performance to AR language models while offering many advantages, such as better speed-quality tradeoffs, controllability, and bi-directional context. Several recent works extend MDMs to multi-modal understanding and generation tasks (Li et al., 2025a; Yu et al., 2025b; Yang et al., 2025; Shi et al., 2025). Compared with the AR+diff setup, unified MDMs avoid the need to carefully tune the balance between AR and diffusion losses by offering a unified objective, resulting in greater simplicity and scalability. Compared with unified AR modeling, unified MDMs offer significantly faster sampling speeds by allowing parallel decoding of multiple tokens.
Despite these advantages, the latest unified MDMs—such as MMaDa (Yang et al., 2025) and Muddit (Shi et al., 2025)—still lag behind state-of-the-art unified AR and AR+diffusion models, both in the breadth of tasks they support and in benchmark performance. There are three main challenges in developing high-performing unified MDMs. First, unified models are expensive to train due to the large size of their language backbones. For example, to build a unified MDM with image generation capability, MMaDa pretrains an 8B model jointly on text and image generation, which is costly. This challenge is further exacerbated by the limited literature on training large-scale masked image generative models. In contrast, many open-sourced large-scale continuous diffusion models such as Flux (Labs, 2024) are readily available. Second, open-sourced resources for masked image generative models (MIGMs) are limited, and the literature on their training techniques and sampling processes is less developed than that for continuous diffusion models. Even the best open-sourced MIGM, Meissonic-1B (Bai et al., 2024), significantly underperforms continuous diffusion models of comparable size (Xie et al., 2025a). Lastly, while these models can perform both understanding and generation tasks, they lack explicit mechanisms to leverage image understanding capabilities to improve generation quality. In fact, MMaDa and Muddit cannot even perform image editing tasks, which require both understanding and generation capabilities. These models simply concatenate text-to-image data and image understanding data during training.
@@
Table 1: Capabilities of different multimodal MDMs. Lavida-O uniquely supports localized understanding, high-resolution image synthesis, image editing and interleaved generation. *Muddit showed examples of simple editing through inpainting. It does not have instruction-based editing capabilities.
@@
LaViDa $\checkmark$ $\times$ $\times$ $\times$ $\times$
Dimple $\checkmark$ $\times$ $\times$ $\times$ $\times$
Muddit $\checkmark$ $\times$ $512^{2}$ * $\times$
@@ 2.1 Masked Diffusion Models
Masked Generative Modeling (MGM) has emerged as an alternative to AR models for modeling sequences of discrete tokens. In MGM, sequences are partially masked during training, and the model learns to predict the masked tokens. Early works such as BERT (Devlin et al., 2019) used MGM as a representation learning objective to extract semantically rich features for downstream tasks. Later works like MaskGIT (Chang et al., 2022) and MUSE (Chang et al., 2023) explored using MGM for generative modeling. In this setup, a sequence is initialized with only mask tokens, which are gradually unmasked by a neural network during sampling to generate the desired output. These works primarily focus on image generation tasks, where discrete image tokenizers such as VQGAN (Esser et al., 2021) are used to convert images into sequences of discrete tokens.
More recently,MDMs (Austin et al., 2021; Sahoo et al., 2024; Lou et al., 2023) have further developed the theory of MGM by formalizing the masking and unmasking process as the forward and reverse diffusion processes in discrete space. This provides a principled framework for training and sampling from these models. MDMs have renewed interest in masked modeling for language generation, offering theoretical advantages over AR models, such as better speed-quality tradeoffs and improved controllability. Notably, LLaDa-8B and Dream-8B (Nie et al., 2025; Ye et al., 2025a) demonstrated that MDMs can achieve competitive performance compared to AR models at scale. Several follow-up works (Li et al., 2025a; Yu et al., 2025b; You et al., 2025; Yang et al., 2025), including LaViDa and MMaDa, extend MDMs to multi-modal tasks such as image understanding and text-to-image generation. Their capabilities are summarized in Table 1.
Formally, given a sequence of $L$ discrete tokens $X_{0}=[X_{0}^{1},X_{0}^{2},\dots,X_{0}^{L}]$, the forward process $q(X_{t}|X_{s})$ gradually masks the tokens over the time interval [0,1], with $1\geq t\geq s\geq 0$. At $t=1$, the sequence $X_{1}=[X_{1}^{1},X_{1}^{2},\dots,X_{1}^{L}]$ consists entirely of masked tokens, denoted by $[M]$. A neural network $p_{\theta}$ is used to model the reverse process $p(X_{s}|X_{t})$. The masked diffusion objective is defined as:
$\mathcal{L}_{\text{MDM}}=-\mathbb{E}_{t,X_{0},X_{t}}\left[\frac{1}{t}\log p_{\theta}(X_{0}|X_{t})\right]=-\mathbb{E}_{t,X_{0},X_{t}}\left[\frac{1}{t}\sum_{i\in\{1,2,\dots,L\},X_{t}^{i}=[M]}\log p_{\theta}(X_{0}^{i}|X_{t})\right]$ (1)
where $\log p_{\theta}(X_{0}|X_{t})$ is factorized into $\prod_{i=1}^{L}p_{\theta}(X_{0}^{i}|X_{t})$ based on independence assumptions (Sahoo et al., 2024). At inference time, the model starts from a fully masked sequence $X_{1}=[M,M,\dots,M]$ and progressively applies the learned reverse process $\log p_{\theta}(X_{0}|X_{t})$ to recover the original tokens. We provide a more detailed formulation of MDMs in Appendix A.1.
@@ 2.2 Unified Multi-modal Models
Unified multi-modal models such as GPT-4o (OpenAI, 2024) are capable of both image understanding and generation tasks, leading to strong performance on tasks requiring both capabilities, such as image editing. Generally, there are two dominant types of unified models based on their modeling objectives. The first type employs an autoregressive next-token prediction objective for text generation and a diffusion objective for image generation (AR+diff). Notable models in this category include BLIP-3o (Chen et al., 2025a), BAGEL (Deng et al., 2025), and Metaqueries (Pan et al., 2025). However, this design involves two different training objectives with distinct numerical scales and training dynamics, often requiring careful tuning of loss weighting and data mixtures.
In contrast, the second type of models employ a unified objectives for both image and texts. Early works like Janus-Pro (Chen et al., 2025c) employ a unified AR modeling objective. Several recent works like MMaDa (Yang et al., 2025) explore a unified masked diffusion objective. Despite some success, a significant performance gap remains between these unified MDMs and state-of-the-art unified models in the AR and AR+diff categories.
Architecturally, unified models also fall into two main categories. The first type, such as Janus and MMaDa, uses a single dense transformer to output both image and text tokens. The second type, such as BAGEL and MetaQueries, employs separate parameter sets for handling image and text modalities. A common design in this category is the mixture-of-transformers (MoT) architecture (Liang et al., 2024), where image and text inputs are processed by different parameter sets but can interact through joint attention mechanisms. These designs are illustrated in Figure 3. While being more flexible, training MoT experts can be expensive due to the duplication of parameters. For example, BAGEL extends a 7B Qwen model into a 7B+7B MoT, doubling the parameter count.
@@ 3.1 Model Architecture
Lavida-O’s model architecture is built on LaViDa (Li et al., 2025a), a diffusion vision-language model capable of image understanding tasks. LaViDa uses a SigLIP (Zhai et al., 2023) vision encoder to convert input images into continuous semantic embeddings $C_{i}$, which are concatenated with token embeddings of text prompts $C_{t}$ to form the final conditional embeddings $C=\text{Concat}(C_{i},C_{p})$ for visual understanding tasks. At each inference step, the diffusion model uses the partially unmasked answer $X_{t}$ and the conditional embedding $C$ to predict the reverse distribution $p_{\theta}(X_{0}|X_{t},C)$, where $X_{0}$ is the clean text answer.
For image understanding tasks, Lavida-O maintains this eact setup of LaViDa. To incorporate visual generation tasks, we extend LaViDa’s design by representing target images as sequences of discrete tokens using a VQ-Encoder (Esser et al., 2021). When performing these tasks, $X_{0}$ and $X_{t}$ contain not only text tokens, but also VQ tokens representing images. For image editing and interleaved generation tasks, we additionally incorporate VQ tokens of input images $C_{v}$ as part of the conditional embedding $C=\text{Concat}(C_{i},C_{v},C_{p})$, since using semantic embeddings $C_{i}$ alone may lose low-level details needed for editing. To reduce the number of tokens and improve computational efficiency, we introduce a token compression module that reduces the number of VQ tokens by a factor of 4. The overall pipeline is illustrated in Figure 2.
@@ 3.1.1 Elastic Mixture-of-Transformers (ElasticMoT)
Our goal is to find an effective and efficient method that can equip LaViDa, an understanding-only diffusion model, with visual generation capabilities. However, both of the existing common choices described in Section 2.2—dense models and MoT—are very expensive. Dense models use the same set of parameters for both understanding and generation tasks, requiring a mix of understanding and generation data during training to prevent catastrophic forgetting, which is not data-efficient. While the MoT setup allows freezing the understanding branch and training only the generation branch for image generation, its architecture doubles the total parameter count, leading to considerable computational overhead. Moreover, given an 8B base understanding model, both setups require training at least 8B parameters for generation tasks from scratch, which is prohibitively expensive.
To address these limitations, we propose Elastic-MoT, a novel architecture design that efficiently adapts an understanding-only model for image generation tasks. Compared with the vanilla MoT architecture, Elastic-MoT introduces two major modifications. First, instead of using equally sized understanding and generation branches, the generation branch has a smaller hidden size. This reduces the parameter count and enables efficient training. We make this design choice based on the observation that many text-to-image models can generate high-quality images with only 2–4B parameters, suggesting that generation tasks may not require as much capacity as understanding tasks (Xie et al., 2025a, b).
Second, given an $N$-layer model, instead of having joint attention at all layers, we only allow text and image modalities to interact in the first $M$ layers. In the remaining K=NMK=N-M layers, text and image tokens interact only within their modality through self-attention. This design activates only partial parameters for different tasks. For example, in Lavida-O’s final design, the generation branch consists of 2.4B new parameters, while the understanding branch has 8B parameters initialized from LaViDa. With $N=32$ layers, we set $M=K=16$. For image generation tasks, we activate all 2.4B parameters in the image generation branch and 4B parameters from the first 16 layers of the understanding branch, leading to 6.4B active parameters. During text-to-image pretraining, we train only the 2.4B parameters in the generation branch, and omit loading the last 16 layers of the understanding branch since they are unused. This enables efficient scaling of text-to-image pretraining without introducing extra understanding data or bearing the large parameter count of MoT. Similarly, the active parameter count is 8B for understanding tasks, and 10.4B for tasks requiring both understanding and generation, such as interleaved generation. The full design of Elastic-MoT is illustrated in Figure 3. We provide additional details and analysis in Appendix A.2 and B.2.
@@ 3.1.2 Modality-aware Masking
One of the challenges in adapting MoT architecture for MDMs is routing—the mechanism to determine which branch should be activated for each token. This is trivial for unified AR MoT models, where the model can simply learn to generate a special token (e.g., [img_start]) to indicate that the next token should use the generation branch. When the last token of an image is generated, the model automatically switches back to the understanding branch to continue text generation. However, this approach is not feasible for MDMs, which decode tokens in parallel and must decide in advance which mask tokens should be routed to the understanding branch and which to the generation branch. A naive solution is to let the user specify the number and location of text and image tokens, but this is difficult for interleaved generation, such as image generation with self-reflection. To address this issue, we design a modality-aware masking process.
Given a sequence of $M$ text tokens and $N$ image VQ tokens, the vanilla forward diffusion process gradually converts it into $M+N$ mask tokens during the time interval $[0,1]$. By contrast, our modality-aware forward process introduces a special timestamp $t_{\text{exp}}\in[0,1]$, at which fully masked image VQ tokens are collapsed into a special $\text{exp}]$ text token. This process is illustrated in Figure 4(a) (Bottom-up). At inference, we assume all mask tokens are text tokens at the beginning. When a $\text{exp}]$ token is generated, we replace it with a sequence of $L_{\text{img}}$ mask tokens, and specify that these tokens will be processed by the generation branch for image synthesis in subsequent forward calls. $L_{\text{img}}$ is determined by a pre-specified image resolution. This process is also illustrated in Figure 4(a) (Top-down). We provide additional details in Appendix A.3.

### v2

# Lavida-O: Elastic Large Masked Diffusion Models for Unified Multimodal Understanding and Generation

@@ Abstract
We propose Lavida-O, a unified Masked Diffusion Model (MDM) for multimodal understanding and generation. Unlike existing multimodal MDMs such as MMaDa and Muddit which only support simple image-level understanding tasks and low-resolution image generation, Lavida-O  presents a single framework that enables image-level understanding, object grounding, image editing, and high-resolution (1024px) text-to-image synthesis. Lavida-O incorporates a novel Elastic Mixture-of-Transformers (Elastic-MoT) architecture that couples a lightweight generation branch with a larger understanding branch, supported by token compression, universal text conditioning and stratified sampling for efficient and high-quality generation. Lavida-O further incorporates planning and iterative self-reflection in image generation and editing tasks, seamlessly boosting generation quality with its understanding capabilities. Lavida-O achieves state-of-the-art performance on a wide range of benchmarks including RefCOCO object grounding, GenEval text-to-image generation, and ImgEdit image editing, outperforming existing autoregressive models and continuous diffusion models such as Qwen2.5-VL and FluxKontext-dev, while offering considerable speedup at inference. These advances establish Lavida-O as a new paradigm for scalable multimodal reasoning and generation.
@@ 1 Introduction
Recently, Masked Diffusion Models (MDMs) (Lou et al., 2023; Sahoo et al., 2024) have emerged as a competitive alternative to AR models. Unlike AR models, MDMs treat token generation as a diffusion process over discrete tokens. In the forward process, the tokens of a sequence are gradually masked. At inference, we start with a sequence of mask tokens and gradually unmask them to obtain a sequence of meaningful tokens. Large-scale experiments in language modeling (Nie et al., 2025; Ye et al., 2025a) show that MDMs can achieve comparable performance to AR language models while offering many advantages, such as better speed-quality tradeoffs, controllability, and bidirectional context. Several recent works extend MDMs to multi-modal understanding and generation tasks (Li et al., 2025a; Yu et al., 2025b; Yang et al., 2025; Shi et al., 2025). Compared with the AR+diff setup, unified MDMs avoid the need to carefully tune the balance between AR and diffusion losses by offering a unified objective, resulting in greater simplicity and scalability. Compared with unified AR modeling, unified MDMs offer significantly faster sampling speeds by allowing parallel decoding of multiple tokens.
Despite these advantages, the latest unified MDMs—such as MMaDa (Yang et al., 2025) and Muddit (Shi et al., 2025)—still lag behind state-of-the-art unified AR and AR+diffusion models, both in the breadth of tasks they support and in benchmark performance. There are three main challenges in developing high-performing unified MDMs. First, unified models are expensive to train due to the large size of their language backbones. For example, to build a unified MDM with image generation capability, MMaDa pretrains an 8B model jointly on text and image generation, which is costly. This challenge is further exacerbated by the limited literature on training large-scale masked image generative models. In contrast, many open-source large-scale continuous diffusion models such as Flux (Labs, 2024) are readily available. Second, open-source resources for masked image generative models (MIGMs) are scarce, and the literature on their training techniques and sampling processes is less developed than that for continuous diffusion models. Even the best open-source MIGM, Meissonic-1B (Bai et al., 2024), significantly underperforms continuous diffusion models of comparable size (Xie et al., 2025a). Lastly, while these models can perform both understanding and generation tasks, they lack explicit mechanisms to leverage image understanding capabilities to improve generation quality. In fact, MMaDa and Muddit cannot even perform image editing tasks, which require both understanding and generation capabilities. These models simply concatenate text-to-image data and image understanding data during training.
@@
Table 1: Capabilities of different multimodal MDMs. Lavida-O uniquely supports localized understanding, high-resolution image synthesis, image editing and interleaved generation.
@@
LaViDa, Dimple, LLaDa-V $\checkmark$ $\times$ $\times$ $\times$ $\times$
Muddit $\checkmark$ $\times$ $512^{2}$ * [^1] Muddit showed examples of simple editing through inpainting. It does not have instruction-based editing capabilities. $\times$
@@ 2.1 Masked Diffusion Models
Masked Generative Modeling (MGM) has emerged as an alternative to AR models for modeling sequences of discrete tokens. Early works such as BERT (Devlin et al., 2019) used MGM as a representation learning objective. Later works (Chang et al., 2022, 2023) such as MaskGIT explored using MGM for generative modeling. In this setup, a sequence is initialized with only mask tokens, which are then gradually unmasked to generate the desired output. In these works, discrete tokenizers like VQGAN (Esser et al., 2021) are used to convert images into discrete tokens.
More recently,MDMs (Austin et al., 2021; Sahoo et al., 2024; Lou et al., 2023) have further developed the theory of MGM by formalizing the masking and unmasking process as the forward and reverse diffusion processes in discrete space. This provides a principled framework for training and sampling from these models. MDMs have renewed interest in masked modeling for language generation, offering theoretical advantages over AR models, such as better speed-quality tradeoffs and improved controllability. Notably, LLaDa-8B and Dream-8B (Nie et al., 2025; Ye et al., 2025a) demonstrated that MDMs can achieve competitive performance compared to AR models at scale. Several follow-up works (Li et al., 2025a; Yu et al., 2025b; You et al., 2025; Yang et al., 2025) such as LaViDa extend MDMs to multi-modal tasks such as image understanding and text-to-image generation. Their capabilities are summarized in Table 1.
Formally, given a sequence of $L$ discrete tokens $X_{0}=[X_{0}^{1},X_{0}^{2},\dots,X_{0}^{L}]$, the forward process $q(X_{t}|X_{s})$ gradually masks the tokens over the time interval [0,1], with $1\geq t\geq s\geq 0$. At $t=1$, the sequence $X_{1}$ consists entirely of masked tokens, denoted by $[M]$. A neural network $p_{\theta}$ is used to model the reverse process $p(X_{s}|X_{t})$. The masked diffusion objective is defined as:
$\mathcal{L}_{\text{MDM}}=-\mathbb{E}_{t,X_{0},X_{t}}\left[\frac{1}{t}\log p_{\theta}(X_{0}|X_{t})\right]$ (1)
where $p_{\theta}(X_{0}|X_{t})$ is factorized into $\prod_{i=1}^{L}p_{\theta}(X_{0}^{i}|X_{t})$ based on independence assumptions (Sahoo et al., 2024). At inference time, the model starts from a fully masked sequence $X_{1}=[M,M,\dots,M]$ and progressively applies the learned reverse process $\log p_{\theta}(X_{0}|X_{t})$ to recover the original tokens. We provide a more detailed formulation of MDMs in Appendix A.1.
@@ 2.2 Unified Multi-modal Models
Unified multi-modal models such as GPT-4o (OpenAI, 2024) are capable of both image understanding and generation tasks, leading to strong performance on tasks requiring both capabilities, such as image editing. Generally, there are two dominant types of unified models based on their modeling objectives. The first type, such as BAGEL (Deng et al., 2025), employs an AR objective for text generation and a diffusion objective for image generation (AR+diff). However, this design involves two different training objectives with distinct numerical scales and training dynamics, often requiring careful tuning of loss weighting and data mixtures. In contrast, the second type of models employ a unified objectives for both image and texts. Early works like Janus-Pro (Chen et al., 2025c) employ a unified AR modeling objective. Recent works like MMaDa (Yang et al., 2025) explore a unified MDM objective. Despite some success, a significant performance gap remains between these unified MDMs and state-of-the-art unified models in the AR and AR+diff categories.
Architecturally, unified models also fall into two main categories. The first type, such as Janus and MMaDa, uses a single dense transformer to output both image and text tokens. The second type, such as BAGEL and MetaQueries, employs separate parameter sets for handling image and text modalities. A common design in this category is the mixture-of-transformers (MoT) architecture (Liang et al., 2024), where image and text inputs are processed by different parameter sets but can interact through joint attention mechanisms. These designs are illustrated in Figure 3. While being more flexible, training MoT experts can be expensive due to their large parameter counts.
@@ 3.1 Model Architecture
Lavida-O’s model architecture is built on LaViDa (Li et al., 2025a), a diffusion model capable of only image understanding tasks. LaViDa uses a SigLIP (Zhai et al., 2023) vision encoder to convert input images into continuous semantic embeddings $C_{i}$, which are concatenated with token embeddings of text prompts $C_{t}$ to form the final conditional embeddings $C=\text{Concat}(C_{i},C_{p})$ for visual understanding tasks. At each inference step, the diffusion model uses the partially unmasked answer $X_{t}$ and the conditional embedding $C$ to predict the clean text answer $X_{0}$.
For image understanding tasks, Lavida-O maintains this exact setup of LaViDa. To incorporate visual generation tasks, we extend LaViDa’s design by representing target images as sequences of discrete tokens using a VQ-Encoder (Esser et al., 2021). When performing these tasks, $X_{0}$ and $X_{t}$ contain not only text tokens, but also VQ tokens that represent images. For image editing and interleaved generation tasks, we additionally incorporate VQ tokens of input images $C_{v}$ as part of the conditional embedding $C=\text{Concat}(C_{i},C_{v},C_{p})$, since using semantic embeddings $C_{i}$ alone can degrade the low-level details needed for editing. To reduce the number of tokens and improve computational efficiency, we introduce a token compression module that reduces the number of VQ tokens by a factor of 4. The overall pipeline is illustrated in Figure 2.
@@ 3.1.1 Elastic Mixture-of-Transformers (ElasticMoT)
Our goal is to find an efficient method that can equip an understanding-only diffusion model with visual generation capabilities. However, both of the existing common choices described in Section 2.2—dense models and MoT—are very expensive. Dense models use the same set of parameters for all tasks, requiring a mix of understanding and generation data during training to prevent catastrophic forgetting, which is not data-efficient. While the MoT setup allows freezing the understanding branch and training only the generation branch for image generation, its architecture doubles the total parameter count, leading to considerable computational overhead. Moreover, given an 8B base understanding model, both setups require training at least 8B parameters for generation tasks from scratch, which is prohibitively expensive.
To address these limitations, we propose Elastic-MoT, a novel architecture design that efficiently adapts an understanding-only model for image generation tasks. Compared with the vanilla MoT architecture, Elastic-MoT introduces two major modifications. First, instead of using equally sized branches, the generation branch has a smaller hidden size. This reduces the parameter count and enables efficient training. We make this design choice based on the observation that many text-to-image models can generate high-quality images with only 2–4B parameters, suggesting that generation tasks may not require as much capacity as understanding tasks (Xie et al., 2025a, b).
Second, given an $N$-layer model, instead of having joint attention at all layers, we only allow text and image modalities to interact in the first $M$ layers. In the remaining K=NMK=N-M layers, text and image tokens interact only within their modality through self-attention. This design activates only partial parameters for different tasks. For example, in Lavida-O’s final design, the generation branch has 2.4B new parameters and the understanding branch 8B parameters from LaViDa. With $N=32$ layers and $M=K=16$, image generation activates only 6.4B parameters (2.4B from generation + 4B from the first 16 understanding layers). During text-to-image pretraining, only the 2.4B generation branch is trainable, further improving the efficiency. Similarly, understanding tasks use 8B active parameters, while interleaved tasks requiring both branches use 10.4B. The full Elastic-MoT design is shown in Figure 3, with further details in Appendix A.2 and B.2.
@@ 3.1.2 Modality-aware Masking
One of the challenges in adapting MoT architecture for MDMs is routing—the mechanism to determine which branch should be activated for each token. This is trivial for unified AR MoT models, where the model can simply learn to generate a special token (e.g., [img_start]) to indicate that the next token should use the generation branch. However, MDMs decode tokens in parallel and must decide in advance which mask tokens should be routed to the understanding branch and which to the generation branch. A naive solution is to let the user specify the number and location of text and image tokens, but this is difficult for interleaved generation, such as image generation with self-reflection. To address this issue, we design a modality-aware masking process.
Given a sequence of $M$ text tokens and $N$ image VQ tokens, the vanilla forward diffusion process gradually converts it into $M+N$ mask tokens during the time interval $[0,1]$. By contrast, our modality-aware forward process introduces a special timestamp $t_{\text{exp}}\in[0,1]$, at which fully masked image VQ tokens are collapsed into a special $\text{exp}]$ text token. This process is illustrated in Figure 4(a) (Bottom-up). At inference, we assume all mask tokens are text tokens at the beginning. When a $\text{exp}]$ token is generated, we replace it with a sequence of $L_{\text{img}}$ mask tokens, and specify that these tokens will be processed by the generation branch for image synthesis in subsequent forward calls. This process is also illustrated in Figure 4(a) (Top-down). We provide additional details in Appendix A.3.
3.2 Task-Specific Designs
Universal Text Conditioning. A common approach to improving the quality of text-to-image models is micro-conditioning (Podell et al., 2023), which conditions the image generation process on extra parameters such as original image resolution, crop coordinates, and image quality scores. This is typically achieved via specialized embeddings. However, since a unified model has strong language understanding and reasoning capabilities, we can simply append these conditions as plain text (e.g., “SCORE: 5.40”) to the end of user prompts. In addition to common conditions, we also incorporate image luminance and contrast as micro-conditions. This simple and effective design not only improves image quality by biasing generation toward high-scoring distributions, but also gives users more refined control over outputs. We provide additional details in Appendix A.4.